# A metrological framework for uncertainty evaluation in machine learning classification models

### v2

@@ Abstract
Machine learning (ML) classification models are increasingly being used in a wide range of applications where it is important that predictions are accompanied by uncertainties, including in climate and earth observation, medical diagnosis and bioaerosol monitoring. The output of an ML classification model is a type of categorical variable known as a nominal property in the International Vocabulary of Metrology (VIM). However, concepts related to uncertainty evaluation for nominal properties are not defined in the VIM, nor is such evaluation addressed by the Guide to the Expression of Uncertainty in Measurement (GUM). In this paper we propose a metrological conceptual uncertainty evaluation framework for ML classification, and illustrate its use in the context of two applications that exemplify the issues and have significant societal impact, namely, climate and earth observation and medical diagnosis. Our framework would enable an extension of the VIM and GUM to uncertainty for nominal properties, which would make both applicable to ML classification models.
@@ 1 Introduction
1.1 Machine learning classification in metrology
The use of machine learning (ML) models offers the potential to bring great benefits to a wide range of societal challenges, for example in climate and earth observation, health care, and advanced manufacturing. ML models can be used to make sense of large data sets and learn complex relationships within them, and also to automate manual processes that require considerable time or expertise.
Many of these applications involve the measurement of some signal on an instrument and inferring information of value to society from that signal. Traditionally, such problems would have involved developing a physical or empirical model that would relate the measurements to the information of interest. Where the information of interest is a quantity or set of quantities, the empirical model developed would form what is considered in metrology a measurement model, providing the relationship between input quantities and a measurand, the “quantity intended to be measured” [JCGMVIM3].
ML methods, and in particular, a type of ML known as supervised learning, where empirical models are built to infer predictions of a specific variable given knowledge of other variables, are increasingly being used to establish (define and parametrise) such models. The phrase “to learn the model” is commonly used in ML to describe model development using supervised learning.
The distinctive feature of supervised learning models is that they are learned from training data, rather than purely using physical intuition or laws. Two types of supervised learning are often distinguished: regression and classification. The distinction is based on the type of variable that is being predicted: a regression model predicts a quantitative variable, while a classification model predicts a categorical variable. In Section 2 we define the terms quantitative and categorical more precisely and provide examples of each.
Both regression and classification models have wide application, but the focus of this paper is on classification models. By a classification model we mean a function $f$, which takes some number of input variables and returns a categorical output variable. Formally speaking, we can denote a collection of input variables by $X$. A collection is like a vector but can include both quantitative and categorical variables. The output categorical variable is denoted by $Y$, and can be one of $K$ classes denoted by $\{c_{1},\dots,c_{K}\}$. Following the convention of the Guide to the Expression of Uncertainty in Measurement (GUM) [JCGMGUM], we describe an instance of these variables, that is the specific implementation given a set of measured or otherwise estimated input variables, using lower case. Thus $x$ is an instance of $X$ and $y$ is an instance of $Y$ and these are related through
@@ 1.1 Machine learning classification in metrology
In reality, most classification models output a particular type of categorical variable known as a nominal variable (sometimes referred to as an unordered categorical variable), or a nominal property in the language of the International Vocabulary of Metrology (VIM) [JCGMVIM3], in which the categories have no natural ordering. See Section 2 for a more detailed discussion of variable types and how they are categorised in the VIM.
More specifically, an ML classification model is one that has been learned from labelled training data, by which we mean $N$ samples (or observations) of paired data $(x_{i},y_{i}),\,i=1,\dots,N$, consisting of observations of both the collection of input variables $X$ and the output variable $Y$. Any implementation of supervised learning consists of two stages: a training phase and a prediction phase. In the training phase, the model $f$ is learned from known paired inputs and outputs, which typically involves optimising the parameters of the model. In the prediction phase, the model is then evaluated upon new input variables.
@@ 1.2 Uncertainty evaluation for machine learning classification
ML classification models are currently being used in several applications in which it is crucial that predictions are accompanied by an evaluation of uncertainty. Examples include earth observation such as land cover classification, mapping of water, ice extent and cloud masking [ESA2022, ESA2023, Skakun_2022], environmental monitoring such as bioaerosol classification [buters_automatic_2024, erb_real-time_2024], and medical diagnosis such as the detection of heart arrythmias from electrocardiogram (ECG) and photoplethysmography (PPG) signals [venton2021robustness, torres2020multi]. See Section 4 where we illustrate the conceptual contributions of this paper in the context of two case studies: land cover classification and atrial fibrillation detection. The importance of uncertainty evaluation for ML was highlighted in the recent Strategic Research Agenda developed by the European Metrology Network for Mathematics and Statistics (MATHMET) [heidenreich_strategic_2023, Section 3].
In a metrology context, the outputs of an ML model may also be used to make inferences about other variables, that is, they can be used as input quantities for further measurement models or classification models. It follows that, as well as being important as an end in itself, uncertainty evaluation also enables the outputs of ML models to be used as part of an ongoing traceability chain, where the term traceability chain is used to include both traditional metrological traceability chains, where artefacts or reference materials are passed between facilities, and situations where data are passed through multiple processing steps (e.g. [woolliams2025metrological]). In Section 3.3, we provide an illustration of how the outputs of an ML classification model, along with their accompanying uncertainties, can be included as inputs to other models.
@@ 1.2.1 The uncertainty evaluation task for machine learning classification models
For classification models, a natural way to evaluate uncertainty is to assign a probability to each of the $K$ possible classes rather than simply returning a single predicted class [JCGMVIM3, Definition 2.9]. Assigning a probability to each class of a nominal property is equivalent to obtaining the probability mass function (PMF) for the classification. In Section 3 we develop the observation that the role of the PMF here is the natural counterpart of the role of the probability density function (PDF) in the measurement of a quantitative variable (quantity). We also show in Section 3 how the PMF of a categorical variable can be used as an input into other models, allowing the propagation of uncertainty as part of the traceability chain.
Section 1.1 described how ML classification is a type of supervised learning where a model is learned that relates outputs to inputs, and where an implementation of ML classification involves a training phase followed by a prediction phase. In ML classification models, contributions to the overall uncertainty of a prediction are introduced in both the training and prediction phases, which is in contrast to rule-based classifiers in which there is no training phase.
@@
    Training data uncertainty. In the training phase, one contribution to prediction uncertainty is measurement uncertainty in the training data itself. For example, when classifying landcover based upon surface reflectance measurements (see our case study in Section 4.1), there is uncertainty in the surface reflectance measurements due to imperfect knowledge of aerosol optical depth (input variables), and the landcover data is also subject to labelling errors (output data). These training data uncertainties affect the training of the model.
@@
    Class assignment uncertainty. In classification tasks for which ML might be used, there is typically no theoretical justification for the choice of input variables, and therefore no reason to assume that observations of the input variables are sufficient to determine a classification unambiguously. Such tasks are in this sense ill-posed: a given observation for the input variables could occur for data drawn from more than one class. In such cases, there is a limitation, inherent in the modelling task, to determining the class to which a data sample belongs. This limitation is itself dependent upon the choice of input variables. Feature selection refers to the common practice in ML of choosing a subset of available input variables, which can help to remove redundant/unnecessary input variables. However, feature selection cannot eliminate class asignment uncertainty if the original selection of variables is insufficient.
@@
    Uncertainty about choice of model. The supervised learning approach relies upon the selection of a certain kind of classification model, and this choice introduces uncertainty. When multiple models are compared, the “best” model is usually selected according to empirical performance, and the reliance upon empirical testing also introduces uncertainty.
@@
    Model parameter uncertainty. ML models need to be optimised using the training data, for example, optimising the values of its parameters. There is, therefore, also uncertainty related to the estimates of the ML model parameters. Some of this uncertainty may be due to the challenges of optimising a model with multiple local optima. The ability to optimise an ML model also depends upon having training data that provides sufficient coverage of the region of data space that is of interest. Training data requirements vary depending upon the complexity of the classification task and the chosen ML model.
@@
Many modern ML classifiers were originally developed with the aim of identifying the (single) correct class. However, the need to go beyond deterministic classifiers and evaluate uncertainty has been recognised in the ML community; see for example [hullermeier2021aleatoric]. Many, though not all, ML classifiers are probabilistic, in the sense that they return class probabilities [murphy2012machine]. Methods differ according to which sources of uncertainty are captured and whether a Bayesian or frequentist approach is adopted.
@@
In our first case study (Section 4.1), a Bayesian approach is taken using a Quadratic Discriminant Analysis (QDA) model [bilson2025uncertaintyawarebayesianmachinelearning]. This method explicitly models four of the five sources of uncertainty described above and illustrated in Figure 1. The method falls into the general category of generative modelling in which the input variables in each class are assumed to follow some statistical distribution, which in the case of QDA is a multivariate Gaussian distribution. Other examples of generative classifiers include Linear Discriminant Analysis and Naïve Bayes; see [murphy2012machine] for more details.
Many modern ML classifiers are instead examples of discriminative modelling in which classification is achieved without making statistical assumptions about the data. In our second case study (Section 4.2), a discriminative modelling approach based on neural network models known as Monte Carlo Dropout is used [gal_dropout_2016, kendall2017uncertainties]. This method captures uncertainty arising from both the data and the model, but it does not explicitly model the various sources of uncertainty in Figure 1. The method does, however, have the advantage of being applicable to high-complexity neural networks, which are known to be especially able to capture complex non-linear relationships. Other examples of discriminative classifiers include logistic regression, support vector machines and tree-based methods such as random forests; see [Hastie2008] for more details.
@@ 1.2.2 Related metrological frameworks
Evaluation of the uncertainty of quantitative outputs from measurement models is addressed by the GUM suite of documents [JCGMGUM, JCGM101, JCGM102] and the related VIM vocabulary [JCGMVIM3]. Central to the GUM is the concept of the propagation of measurement uncertainty through measurement models whose parameters have been determined. This scenario would correspond to a trained ML model in which estimates of the parameters are used. The GUM can also treat the parameters that define the measurement model as input quantities in their own right, propagating uncertainty (and where there are multiple such parameters, their covariance) as with other input quantities.
The GUM framework is explained for models obtained through regression of parametric models (empirical measurement models) in GUM-6 [jcgm-2020gum6]. Uncertainty associated with model parameter estimates is addressed in [jcgm-2020gum6, clause 7]. Uncertainty concerning the choice of model is discussed briefly in [jcgm-2020gum6, clause 8.5]. The inclusion of effects due to measurement uncertainties in either the input or output variables is addressed in [jcgm-2020gum6, clauses 9 and 10].
The GUM framework can thus, in principle, be applied to any ML regression model. However, the GUM documents mentioned above do not address classification models since the concept of measurement uncertainty evaluation has been developed in the metrology community (and in the GUM and VIM documents) for quantitative variables, or quantities in the language of the VIM and the GUM. The output of a classification model, on the other hand, is not a quantitative variable, but rather an unordered categorical variable, or a nominal property.
A certain kind of classification model is addressed in the fourth document of the GUM suite dedicated to conformity assessment [JCGM106-2012]. In conformity assessment, the result of a measurement, which represents a quantity (e.g., mass concentration of mercury in a sample of industrial water), is converted into a nominal variable, which has two categories: conforming and non-conforming. The conversion is implemented by identifying whether the result of a measurement falls into an interval of permissible values set by one or two tolerance limits. Such conversion essentially represents a binary classification task. Binary classification performed by applying a threshold to a quantity (similar to conformity testing) has many applications. For example, in earth observation, operational cloud masking of satellite imagery has been historically performed by applying empirically derived thresholds to top-of-atmosphere reflectance measurements, e.g., see [Louis2021] (Figure 2). However, it is important to note that with the recent progress of ML learning, and specifically deep learning, more efforts have been made in developing more advanced cloud masking algorithms (e.g, [LopezPuigdollers2021]).
@@
Similarly to a probabilistic ML classifier, in conformity assessment the assigned category is provided with a probability value, the conformance probability. However, in conformity assessment the classifier is rule-based, there is no training phase, and therefore uncertainty related to a training phase does not need to be taken into account. For these reasons, the framework in [JCGM106-2012] cannot be used for a comprehensive uncertainty evaluation of ML classifiers, which rely on training.
In a standard conformity assessment a threshold is fixed (e.g., a certain critical mass concentration of mercury based on which a water sample is considered to be conforming or non-conforming). In contrast, in a binary classification performed alike conformity assessment (e.g, cloud masking described above), a threshold is often empirically derived (e.g., based on expert knowledge of top-of-atmosphere reflectance of a cloud vs. non-cloud). In the latter case, there would be an additional uncertainty source associated with the choice of the threshold, which will affect the classification results, and this uncertainty source is not considered in a standard conformity assessment [JCGM106-2012].

### v3

@@ Abstract
Machine learning (ML) classification models are increasingly being used in a wide range of applications where it is important that predictions are accompanied by uncertainties, including in climate and earth observation, medical diagnosis and bioaerosol monitoring. The output of an ML classification model is a type of categorical variable known as a nominal property in the International Vocabulary of Metrology (VIM). However, concepts related to uncertainty evaluation for nominal properties are not defined in the VIM, nor is such evaluation addressed by the Guide to the Expression of Uncertainty in Measurement (GUM). In this paper we propose a metrological conceptual uncertainty evaluation framework for nominal properties. This framework is based on probability mass functions and summary statistics thereof, and it is applicable to ML classification. We also illustrate its use in the context of two applications that exemplify the issues and have significant societal impact, namely, climate and earth observation and medical diagnosis. Our framework would enable an extension of the GUM to uncertainty for nominal properties, which would make both applicable to ML classification models.
@@ 1 Introduction
1.1 Machine learning classification from a metrology perspective
The use of machine learning (ML) models offers the potential to bring great benefits to a wide range of societal challenges, for example in climate and earth observation [1, 2, 3, 4, 5], health care [6, 7], and advanced manufacturing [8]. ML models can be used to make sense of large data sets and learn complex relationships within them, and also to automate manual processes that require considerable time or expertise.
Many of these applications involve the measurement of some signal on an instrument and inferring information of value to society from that signal. Traditionally, such problems would have involved developing a physical or empirical model that would relate the measurements to the information of interest. Where the information of interest is a quantity or set of quantities, the empirical model developed would form what is considered in metrology a measurement model, providing the relationship between input quantities and a measurand, the “quantity intended to be measured” [9].
ML methods, and in particular methods of supervised learning, are increasingly being used to establish (define and parametrise) measurement models. In supervised learning, data-driven models are built to infer predictions of the value of some (output) variable given knowledge of the values of other (input) variables. Typically a model type is first selected, and then its parameters are optimised. An example of a type of model used in supervised learning is a neural network with a predefined architecture (see Section 4.2). The phrase “to learn the model” is commonly used in ML to refer to the optimisation of the model parameters.
The distinctive feature of supervised learning models is that their parameters are learned from training data, rather than purely using physical intuition or laws. Two types of supervised learning are often distinguished: regression and classification. The distinction is based on the type of variable that is being predicted: a regression model predicts a quantitative variable, while a classification model predicts a categorical variable. In Section 2 we define the terms quantitative and categorical more precisely and provide examples of each.
Both regression and classification models have wide application, but the focus of this paper is on classification models. By a classification model we mean a function $f$, which takes some number of input variables and returns a categorical output variable. Formally speaking, we can denote a collection of input variables by $X$. A collection is like a vector but can include both quantitative and categorical variables. The output categorical variable is denoted by $Y$, and can take one of $K$ values (often referred to in ML as classes) denoted by $\{c_{1},\dots,c_{K}\}$. Following the convention of the Guide to the Expression of Uncertainty in Measurement (GUM) [10], we denote a value taken by any of these variables using lower case. Thus $x$ is an instance of $X$ and $y$ is an instance of $Y$ and these are related through
@@ 1.1 Machine learning classification from a metrology perspective
In reality, most classification models output a particular type of categorical variable known as a nominal variable (sometimes referred to as an unordered categorical variable), or a nominal property in the language of the International Vocabulary of Metrology (VIM) [9], in which the categories have no natural ordering. See Section 2 for a more detailed discussion of variable types and how they are categorised in the VIM.
More specifically, an ML classification model is one that has been learned from labelled training data, by which we mean $N$ samples (or observations) of paired data $(x_{i},y_{i}),\,i=1,\dots,N$, consisting of observations of both the collection of input variables $X$ and the output variable $Y$. The supervised learning process consists of two stages: a training phase and a prediction phase. In the training phase, the model $f$ is learned from known paired inputs and outputs, which typically involves optimising the parameters of the model. In the prediction phase, the model is then evaluated upon new input variables. ML classification models, which are learned from training data, can be distinguished from rule-based classifiers which use predetermined classification rules which are not dependent upon data.
@@ 1.2 Uncertainty evaluation for machine learning classification
It is always important that an ML classification model is deployed in a trustworthy and quality-assured way. More specifically, ML classification models are currently being used in several applications in which it is crucial that predictions are accompanied by an evaluation of uncertainty. Examples include earth observation such as land cover classification, mapping of water, ice extent and cloud masking [1, 2, 3], environmental monitoring such as bioaerosol classification [4, 5], and medical diagnosis such as the detection of heart arrythmias from electrocardiogram (ECG) and photoplethysmography (PPG) signals [6, 7]. See Section 4 where we illustrate the conceptual contributions of this paper in the context of two case studies: land cover classification and atrial fibrillation detection. The importance of uncertainty evaluation for ML was highlighted in the Strategic Research Agenda developed by the European Metrology Network for Mathematics and Statistics (MATHMET) [11, Section 3].
In a metrology context, the outputs of an ML model may also be used to make inferences about other variables, that is, they can be used as input quantities for further measurement models or classification models. It follows that, as well as being important as an end in itself, uncertainty evaluation also enables the outputs of ML models to be used as part of an ongoing traceability chain, where the term traceability chain is used to include both traditional metrological traceability chains, where artefacts or reference materials are passed between facilities, and situations where data are passed through multiple processing steps (e.g. [12]). In Section 3.3, we provide an example of how the outputs of an ML classification model, along with their accompanying uncertainties, can be included as inputs to other models.
@@ 1.2.1 The uncertainty evaluation task for machine learning classification models
For classification models, a common way to express the confidence of a prediction is to assign a probability to each of the $K$ possible classes rather than simply returning a single predicted class [9, Definition 2.9]. Assigning a probability to each class of a nominal property is equivalent to obtaining the probability mass function (PMF) for the classification. In Section 3 we develop the observation that the role of the PMF here is the natural counterpart of the role of the probability density function (PDF) in the measurement of a continuous quantitative variable (continuous quantity). We also show in Section 3 how the PMF of a categorical variable can be used as an input into other models, allowing the propagation of uncertainty through multistage measurement models involving nominal properties.
Section 1.1 described how ML classification is a type of supervised learning where a model is learned that relates outputs to inputs, and where an implementation of ML classification involves a training phase followed by a prediction phase. In ML classification models, contributions to the overall uncertainty of a prediction are introduced in both the training and prediction phases, which is in contrast to rule-based classifiers in which there is no training phase and where the classification rule is predetermined, for example tolerance thresholds in conformity assessment (see also Section 1.2.2).
A crucial assumption in supervised learning is that the model learned on training data is applicable to input data for which predictions are desired. One way to statistically characterise this assumption is to say that the input data (for both training and prediction) are independent and identically distributed (i.i.d.). In practice, this assumption must be validated through appropriate experimentation. For example the i.i.d. assumption can be validated using a suitable hypothesis test. Where the assumption is violated, new input data for which model predictions are required is said to have undergone distribution shift; see for example [13, Chapter 8]. Uncertainty evaluation methods for supervised learning predictions typically assume the absence of distribution shift, and such methods are only reliable to the extent that this assumption is not violated.
@@
    Training data uncertainty. In the training phase, one contribution to prediction uncertainty is uncertainty in the training data itself. For example, when classifying landcover based upon surface reflectance measurements (see our case study in Section 4.1), there is uncertainty in the surface reflectance measurements due to imperfect knowledge of aerosol optical depth (input variables), and the landcover data is also subject to labelling errors (output data). These training data uncertainties affect the training of the model.
@@
    Class assignment uncertainty. In classification tasks for which ML might be used, there is typically no theoretical justification for the choice of input variables, and therefore no reason to assume that observations of the input variables are sufficient to determine a classification unambiguously. Such tasks are in this sense ill-posed: a given observation for the input variables could occur for data drawn from more than one class. In such cases, there is a limitation, inherent in the modelling task, to determining the class to which a data sample belongs. This limitation is itself dependent upon the choice of input variables. Feature selection refers to the common practice in ML of choosing a subset of available input variables, which can help to remove redundant/unnecessary input variables. However, feature selection cannot eliminate class assignment uncertainty if the original selection of variables is insufficient.
@@
    Uncertainty about choice of model. The supervised learning approach relies upon the selection of a certain type of classification model (see Section 1.1). The model type is typically selected using empirical testing, and this process introduces uncertainty. Bayesian model averaging is one method that has been proposed to capture uncertainty due to choice of model [14].
@@
    Model parameter uncertainty. ML models need to be optimised using the training data, for example, optimising the values of its parameters. There is, therefore, also uncertainty related to the estimates of the ML model parameters. Some of this uncertainty may be due to the challenges of optimising a model with multiple local optima. The ability to optimise an ML model also depends upon having training data that samples the region of data space that is of interest sufficiently densely. Training data requirements vary depending upon the complexity of the classification task and the chosen ML model.
@@
Many modern ML classifiers were originally developed with the aim of identifying the (single) correct class. However, the need to go beyond deterministic classifiers and evaluate uncertainty has been recognised in the ML community; see for example [15]. Many, though not all, ML classifiers are probabilistic, in the sense that they return class probabilities [16]. Methods differ according to which sources of uncertainty are captured and whether a Bayesian or frequentist approach is adopted.
@@
In our first case study (Section 4.1), a Bayesian approach is taken using a Quadratic Discriminant Analysis (QDA) model [17]. This method explicitly models four of the five sources of uncertainty described above and illustrated in Figure 1. The method falls into the general category of generative modelling in which the input variables in each class are assumed to follow some statistical distribution, which in the case of QDA is a multivariate Gaussian distribution. Other examples of generative classifiers include Linear Discriminant Analysis and Naïve Bayes; see [16] for more details.
Many modern ML classifiers are instead examples of discriminative modelling in which classification is achieved without making statistical assumptions about the data. In our second case study (Section 4.2), a discriminative modelling approach based on neural network models known as Monte Carlo Dropout is used [18, 19]. This method captures uncertainty arising from both the data and the model, but it does not explicitly model the various sources of uncertainty in Figure 1. The method does, however, have the advantage of being applicable to high-complexity neural networks, which are known to be especially able to capture complex non-linear relationships. Other examples of discriminative classifiers include logistic regression, support vector machines and tree-based methods such as random forests; see [20] for more details.
@@ 1.2.2 Related metrological frameworks
Evaluation of the uncertainty of quantitative outputs from measurement models is addressed by the GUM suite of documents [10, 21, 22, 23] and the related VIM vocabulary [9]. Central to the GUM is the concept of the propagation of measurement uncertainty through measurement models whose parameters have been determined. This scenario would correspond to a trained ML model in which estimates of the parameters are used. The GUM can also treat the parameters that define the measurement model as input quantities in their own right, propagating uncertainty (and where there are multiple such parameters, their covariance) as with other input quantities.
The GUM framework is explained for models obtained through regression of parametric models (empirical measurement models) in GUM-6 [24]. Uncertainty associated with model parameter estimates is addressed in [24, clause 7]. Uncertainty concerning the choice of model is discussed briefly in [24, clause 8.5]. The inclusion of effects due to measurement uncertainties in either the input or output variables is addressed in [24, clauses 9 and 10].
The frameworks developed in the GUM suite of documents can thus, in principle, be applied to any ML regression model. However, the GUM documents mentioned above do not address classification models since the concept of measurement uncertainty evaluation has been developed in the metrology community (and in the GUM and VIM documents) for quantitative variables, or quantities in the language of the VIM and the GUM. The output of a classification model, on the other hand, is not a quantitative variable, but rather an unordered categorical variable, or a nominal property.
A certain kind of classification model is addressed in the fourth document of the GUM suite dedicated to conformity assessment [25]. In conformity assessment, the result of a measurement, which represents a quantity (e.g., mass concentration of mercury in a sample of industrial water), is converted into a nominal variable, which has two categories: conforming and non-conforming. The conversion is implemented by identifying whether the result of a measurement falls into an interval of permissible values set by one or two tolerance limits. Such conversion essentially represents a binary classification task. Binary classification performed by applying a threshold to a quantity (similar to conformity testing) has many applications. For example, in earth observation, operational cloud masking of satellite imagery has been historically performed by applying empirically derived thresholds to top-of-atmosphere reflectance measurements, e.g., see [26] (Figure 2). However, it is important to note that with the recent progress of ML learning, and specifically deep learning, more efforts have been made in developing more advanced cloud masking algorithms (e.g, [27]).
@@
Similarly to a probabilistic ML classifier, in conformity assessment the assigned category is provided with a probability value, the conformance probability. However, in conformity assessment the classifier is rule-based, there is no training phase, and therefore uncertainty related to a training phase does not need to be taken into account. For these reasons, the framework in [25] cannot be used for a comprehensive uncertainty evaluation of ML classifiers, which rely on training.
In a standard conformity assessment a threshold is fixed (e.g., a certain critical mass concentration of mercury based on which a water sample is considered to be conforming or non-conforming). In contrast, in a binary classification performed alike conformity assessment (e.g, cloud masking described above), a threshold is often empirically derived (e.g., based on expert knowledge of top-of-atmosphere reflectance of a cloud vs. non-cloud). In the latter case, there would be an additional uncertainty source associated with the choice of the threshold, which will affect the classification results, and this uncertainty source is not considered in a standard conformity assessment [25].